jack and jerry can complete a job together in 4 hours. if jack require 6 hours to do the job alone, how many hours will it take jerry to do the job alone.
A.12
B.10
C.8
D.6
E.4

Answer by ptaylor(2198)

Let x=time it takes Jerry to do the job alone
Then Jerry works at the rate of 1/x of the job per hour
And Jack works at the rate of 1/6 of the job per hour
Together they work at the rate of 1/4 of the job per hour
Sooooo
1/x + 1/6= 1/4 multiply each term by 12x to get rid of fractions
12 + 2x=3x subtract 2x from each side
x=12 hrs
ck
1/12 + 1/6 =? 1/4
1/12 + 2/12= 1/4
1/4=1/4
Hope this helps----ptaylor